OpenStudy (anonymous):

So I'm looking for some help on Calculus A, chapter 2 on differentiation. I know how to do the gist of it but i have difficulty with the chain rule. y=x²sine³(2x)

OpenStudy (anonymous):

Let's do this step by step?

OpenStudy (anonymous):

I know that in the chain rule you use \[f \prime( g(x))\times g \prime(x)\]

OpenStudy (anonymous):

i know the following... well i figured it out: f(x)=x²(u) g(x)=sin³(x) f’(x)= 3u² g’(x)=3cos²x

OpenStudy (anonymous):

That is correct But you see here we have a chain rule and a product rule

OpenStudy (anonymous):

well you do all the chain rule of sin³(2x) and then you just do product rule with x² and whatever you got with the product rule right??

OpenStudy (anonymous):

Yes exactly, do that now and ill check if you're right

OpenStudy (anonymous):

okay so i have so far that the chain rule thing is \[3(\sin ^{3}(x))^{2}(3cosx)\] which simplifies to \[(3\sin ^{5} x)(3\cos²x)\] and then \[9\sin ^{5} xcos²x\]

OpenStudy (anonymous):

and now is where i stop because i am lost and don't know how to do the rest.... in fact I'm not even sure i did that part right

OpenStudy (anonymous):

That is incorrect The chain rule is as follows \[\frac{ d }{dx } (\sin(2x))^3\]

OpenStudy (anonymous):

And then....

OpenStudy (anonymous):

hommie where did you learn your chain rule?!?!

OpenStudy (anonymous):

thats implicit differentiation not the chain rule!!

OpenStudy (anonymous):

\[3(\sin(2x))^3 \times 2(\sin(2x)) \times \cos(2x)\]

OpenStudy (anonymous):

never mind i get it

OpenStudy (anonymous):

actually, no this is the chain rule Implicit is taking the derivative of y too.

OpenStudy (anonymous):

why is it that and not f '(g(x))(g '(x)

OpenStudy (anonymous):

because you just rearranged the original equation and got a whole bunch of stuff out of it

OpenStudy (anonymous):

Well no you have repeated chain rule for this case. I didnt rearrange the equation I just took the derivative for the chain rule term You're still not done You gotta do the product rule

OpenStudy (anonymous):

like i get how you got all the numbers but how do you know when to simplify it the way you did and when to do it the way i do?

OpenStudy (anonymous):

i know.... let me finish the problem on paper and get back to you in a bit.

OpenStudy (anonymous):

Well you look at whether you took the derivative of every function within the chain of functions you have to know your f(x) and g(x) when it comes to f(x) and g(x) Btw I did it using the actual way which is the "way you do" Like i said, you have repeated chain rule.

OpenStudy (anonymous):

are you absolutely sure your solution is correct? or am i just going crazy and doing it completely wrong?

OpenStudy (anonymous):

Proof^

OpenStudy (anonymous):

what the heck like literally everything i do is not that answer.....

OpenStudy (anonymous):

ohh goodness i thought i was going crazy..... \[2(3\cos ^{2}(x)\sin ^{3}(2x)\]

OpenStudy (anonymous):

Not it

OpenStudy (anonymous):

i know this

OpenStudy (anonymous):

I would explain but I'm pretty tired

OpenStudy (anonymous):

well i don't get it but its fine ill figure it out thanks for the help.